OpenStudy (anonymous):

The length of the hypotenuse of an isosceles right triangle is sqrt6 inches. How long, in inches, are the legs of the triangle?

OpenStudy (anonymous):

use this sin90=perp/hypt =>perp=hypt*sin90 =>perp=sqrt6 *1=sqrt6 inches

OpenStudy (anonymous):

\[\sqrt{x ^{2}+x ^{2}}=\sqrt{2x ^{2}}=\sqrt{6}\] and solve for x (x is the side length

OpenStudy (raden):

 let x is the length of legs of the triangle. x sqrt(2) = sqrt(6) solve for x

OpenStudy (anonymous):

@Catch.me could u help me with one more?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

np